# Alejandro Benito-Santos, Adrián Ghajari Espinosa, Victor Fresno UNED research center in Natural Language Processing and Information Retrieval

### v1

@@ Abstract
In this article we present UNED-ACCESS 2024, a bilingual dataset that consists of 1003 multiple-choice questions of university entrance level exams in Spanish and English. Questions are originally formulated in Spanish and translated manually into English, and have not ever been publicly released. A selection of current open-source and proprietary models are evaluated in a uniform zero-shot experimental setting both on the UNED-ACCESS 2024 dataset and on an equivalent subset of MMLU questions. Results show that (i) reasoning questions are challenging for models, (ii) smaller models perform worse than larger models and degrade faster in Spanish than in English and (iii) the performance gap between languages is negligible for the best models and grows up to 37% for smaller models. Model ranking on UNED-ACCESS 2024 is almost identical in English and Spanish, and has also a high correlation (0.98 Pearson) with ranking on MMLU, suggesting that a small dataset is sufficiently diverse and representative to measure performance by discipline.
@@
Eva Sánchez Salido, Roser Morante, Julio Gonzalo, Guillermo Marco, Jorge Carrillo-de-Albornoz, Laura Plaza, Enrique Amigó, Andrés Fernández, Alejandro Benito-Santos, Adrián Ghajari Espinosa, Victor Fresno UNED research center in Natural Language Processing and Information Retrievalthanks: nlp.uned.es Correspondence: julio@lsi.uned.es
@@ 1 Introduction
With the recent progress in broadening the generalisation capabilities of Large Language Models (LLMs), much current research focuses on understanding their capabilities and limitations. Evaluations of the most recent generative models, such as Llama-2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024a), Gemini (Anil et al., 2024), Gemma (Mesnard et al., 2024), GPT-3.5 (Brown et al., 2020), GPT-4 and GPT-4o  (Achiam et al., 2024), attempt at measuring their world knowledge, memorisation and inference capabilities through benchmarks designed for these purposes. Although many benchmarks have been proposed as single-task evaluations, with the emergence of general language models such as BERT (Devlin et al., 2019), the development of more comprehensive benchmarks to measure the general capabilities of these models in a multi-task setting has become popular (GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2020), Big-Bench (Srivastava et al., 2022), Big-Bench Hard (Suzgun et al., 2022), MMLU (Hendrycks et al., 2021a), HELM (Liang et al., 2023)). Since most of these benchmarks propose assessments that are performed on artificially created datasets for specific NLP tasks, emphasis is now put on the importance of assessments focusing on human-level cognitive tasks in real-world scenarios, such as exams. In this context, multiple-choice questions have emerged as one of the preferred methods for evaluating new generative models with datasets such as RACE (Lai et al., 2017) or AGIEval (Zhong et al., 2023), which includes university entrance exams, law school entrance exams, mathematics competitions and lawyer qualification tests.
In order to contribute to the assessment of language models within real scenarios, we present a new bilingual dataset of private exam questions, UNED-ACCESS 2024. The dataset contains 1003 multiple-choice questions in Spanish from various subjects of the UNED Access Course for Over-25s in Spanish, and high-quality English professional translations. Two characteristics make this dataset unique: first, the contamination level should be minimal because UNED typically does not release the answers to the exam questions, which are only accessible to the teachers of each course. Second, this is a high-quality bilingual dataset, with original questions in Spanish translated into English manually by a professional translators who did not use any external software.
@@
Our main contributions are: (i) The creation of a new, private benchmark with minimal contamination for the evaluation of generative LLMs on entrance university exams in Spanish (native) and English (translated); (ii) A systematic evaluation of LLMs with the same prompts and hyperparameters for all models, in English and Spanish. (iii) Empirical evidence that the effectiveness gap of LLMs between English and Spanish is inversely correlated with model quality and size: smaller models have lower performance values in both languages and also a larger gap between their effectiveness in English and Spanish. (iv) Empirical evidence that the relative performance of models on UNED-ACCESS 2024 is almost identical to their performance on an equivalent MMLU dataset and that results correlate with the public MMLU results, which may imply that a small sufficiently diverse dataset is representative to measure model performance.
@@ 2 Related work
Standard question/answer datasets include tasks such as common sense reasoning: HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), ARC (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), CommonsenseQA (Talmor et al., 2019), PIQA (Bisk et al., 2019), SIQA (Sap et al., 2019); world knowledge: TriviaQA (Joshi et al., 2017), NaturalQuestions (Kwiatkowski et al., 2019); reading comprehension: RACE (Lai et al., 2017), QuAC (Choi et al., 2018) and SQuAC (Rajpurkar et al., 2016); mathematical reasoning: MATH (Hendrycks et al., 2021b), GSM8K (Cobbe et al., 2021)); and code generation: HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021). Aggregated benchmarks like MMLU (Hendrycks et al., 2021a), AGIEval (Zhong et al., 2023), Big-Bench (Srivastava et al., 2022), and Big-Bench Hard (Suzgun et al., 2022) are also commonly used. In particular, since the introduction of GPT-3, evaluations that comprise a variety of datasets have been standardised (see for instance evaluations of GPT-3 (Brown et al., 2020), GPT-4 (Achiam et al., 2024), Llama and Llama-2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023) or Claude-3 (Anthropic, 2024)).
@@
Recent research has increasingly focused on multilingual evaluations of models like ChatGPT. For instance, Lai et al. (2023) evaluated ChatGPT on 37 languages ranging from high to extremely low resources, across seven tasks in a zero-shot learning setting. Their findings, along with Bang et al. (2023), show that models struggle to generalize in low-resource languages. Similarly, M3Exam (Zhang et al., 2023) assesses models like GPT-4 across nine languages, revealing persistent difficulties in handling low-resource and non-Latin script languages. GPT-4’s technical report itself extends its evaluation to a multilingual version of MMLU self-translated via Azure Translate (Achiam et al., 2024).
@@ 4 Experimental setup
Experiments were conducted using 12 generative models: four proprietary models (GPT-4-Turbo (Achiam et al., 2024), GPT-4o, GPT-3.5-Turbo (Brown et al., 2020) and Claude-3-Opus (Anthropic, 2024)) and eight open-source models (Llama-2-7B (Touvron et al., 2023a), Llama-3-8B (Meta, 2024), Llama-3-70B, Gemma-7B (Mesnard et al., 2024), Gemma-2-27B, Mixtral-8x7B (Jiang et al., 2024a), Mistral-7B (Jiang et al., 2023) and Leniachat-Gemma-2B), all trained for instruction following. In Appendix B.1 details are provided about how the models were accessed.
We decided to apply a uniform zero-shot setting because it more closely resembles the real-world scenario in which people interact with LLMs, and it also results in a simpler and more replicable experimentation. As for the hyperparameters, the temperature was set to 0 in all models to ensure deterministic responses, minimizing creativity and focusing on factual or reasoning-based answers, which suit the dataset. The questions were provided one at a time and the prompt is fixed, providing a system prompt, a user prompt and an assistant prompt. As extra information, the prompt provided the name of the subject to which the question belongs. The language of the prompt was always in the same language of the dataset used to evaluate the model (English or Spanish) as in Zhang et al. (2023) and Achiam et al. (2024). The prompts are the following for Spanish (ES) and English (EN):
@@
In the literature, evaluation is typically based on accuracy, the proportion of correct answers ($C$) over the total responses ($N$). However, to allow for global comparisons across subjects with varying numbers of answer choices, we use Cohen’s Kappa coefficient, which accounts for the number of possible choices ($M$) in multiple-choice questions:
@@ 5 Results
Llama-3-70B-Instruct open-source 0.65 0.67 -3.44
Gemma-2-27B-Instruct open-source 0.64 0.66 -3.40
@@
Mixtral-8x7B-Instruct open-source 0.56 0.57 -2.32
Llama-3-8B-Instruct open-source 0.51 0.50 2.73
Mistral-7B-Instruct open-source 0.46 0.43 5.67
Gemma-7B-It open-source 0.41 0.38 8.93
Llama-2-7B-Chat open-source 0.32 0.25 26.04
Leniachat-Gemma-2B open-source 0.15 0.11 37.27
@@
Table 3 shows the average Cohen’s Kappa results by model in Spanish and English. Additional detailed results per language, model and subject are provided in Table 5 in Appendix C.1. The best overall results for Spanish are achieved by the proprietary models Claude-3, GPT-4-Turbo and GPT-4o, followed by the open models Llama-3-70B, Gemma-2-27B and Mixtral-8x7B which outperform GPT-3.5-Turbo. In the pass margin (around 0.50 points) there is Llama-3-8B, and below are Mistral-7B, Gemma-7B, Llama-2-7B and Leniachat-2B. For English, the ranking of the models by average performance is almost identical to that of Spanish, except for an exchange in the positions of GPT-3.5 and Mixtral, (GPT-3.5 performs better than Mixtral in English but worse in Spanish). At the quantitative level, the results are also very similar in English and Spanish.
@@
Figure 1 shows the average results for each model in Spanish and English, sorted by the mean in Spanish. Five out of twelve models perform slightly better in Spanish than English, two of them proprietary (which perform the best in Spanish overall) and three open (also the three best models among the open-source ones). This does not imply that these models are better in Spanish: English questions are manual translation of the original Spanish ones, and there are at least three language-related factors that affect overall performance: (i) even with high-quality manual translations, there might be translation artifacts (e.g. unusual translation choices) that make the English text harder to process; (ii) contamination: even if the questions are private, the chance that some questions have been seen by the models in Spanish is higher than in English; (iii) language competence of the models in English vs Spanish. Presumably, in the models that perform better in Spanish, effects (i) and (ii) out-weight pure language competence.
Figure 1: Performance of each model on UNED-ACCESS 2024 per language averaged across all subjects.
Among the proprietary models, three of them have similar performance differences (in absolute value), whilst that of GPT-3.5 is the most pronounced, in favour of English. As for the open models, the two best performing ones (Llama-3-70B and Gemma-2) have a very similar performance difference, both in favour of Spanish (and larger that the ones observed for the best proprietary models). Mixtral is the next model in the ranking and still favours Spanish, while all the other models, which do not pass (or are at the limit), favour English. In this set of models the performance difference is very pronounced, specially for Llama-2 and Leniachat, which are the worst performing models. In summary, for the best models we observe a slight advantage for Spanish, while in the worst models the advantage leans towards English. The best models are in turn larger, and therefore more likely to have been trained with more Spanish data, and more data from the web in general, whereas smaller models may rely more solely on English. However, in principle this is not the case for Leniachat, which has been conceived specifically for Spanish. The size of the models is crucial: although we do not know for certain the size of the proprietary models, in general model size is the best performance predictor, with one exception: the very similar performance between Llama-3-70B and Gemma-2-27B, with such a large difference in size, is noteworthy.
We calculate the effectiveness gap between languages from the averaged results as:
@@
Table 3, presented above, shows the percentage effectiveness gap. For Claude-3, GPT-4, Llama-3-70B, Gemma-2-27B and Mixtral-8x7B the gap is negative, which means that they perform better in Spanish. The rest of the models are better in English. There are two cases where the gap is extremely pronounced (in favour of English), coinciding with the two worst-performing models: Leniachat, which is the smallest model, and Llama-2, which is the oldest of the evaluated models.
There is a strong, statistically significant inverse correlation between model quality and language gap: Pearson’s correlation between English performance and EN-ES gap is -0.87 with a p-value of 0.0002; and Pearson’s correlation between performance in Spanish and language gap is -0.89 (p=9.83e-05). The performance of less powerful models degrades much faster in Spanish than in English.
@@
A direct comparison between the public results on MMLU (Hendrycks et al., 2021a) and UNED-ACCESS 2024 is not adequate: the dataset sizes are different (the test split of MMLU has more than 14000 instances) and the subjects and levels differ: MMLU ranges in difficulty from an elementary level to an advanced professional level, whereas UNED-ACCESS 2024 covers only university-entrance level. Also, most public evaluations of MMLU are 5-shot and even include chain of thought prompting, as compared to the zero-shot setting of our experiments. For these reasons we have performed our own evaluation of the models, using the same methodology described earlier, and selecting the subset of 2822 MMLU questions (MMLU-10) from ten topics which correspond most closely to the level and topics of UNED-ACCESS: high school biology, high school chemistry, high school computer science, elementary mathematics, high school mathematics, high school physics, high school macroeconomics, high school microeconomics, high school psychology and high school world history.
@@
First, we have checked whether the results that we obtain on MMLU-10 with our methodology correlate with the published results (details are provided in Appendix C.3), obtaining a positive answer. This allows us to compare the results obtained on MMLU-10 with the results on UNED-ACCESS 2024 in English and check whether they correlate. Table 4 shows both results in terms of Cohen’s Kappa (recall that the MMLU questions all have 4 answers). Pearson’s correlation between these results is 0.98 (p=5.04e-08): both results have almost perfect, statistically significant correlation. This consistency reinforces the reliability of results obtained with both datasets.
The fact that the results on both datasets are so similar, in spite of the fact that UNED-ACCESS should have minimal contamination, implies that either the questions in UNED-ACCESS are somewhat easier in average, or that there is some kind of indirect contamination: some questions might be recurring for some subjects, and models might have seen variations of the questions even if they do not have access to the original ones. Also, the similarity of the results seems to indicate that the translation effect is not particularly relevant.
In Table 6 of Appendix C.2 we also provide the evaluation results on the MMLU-10 dataset per subject, which shows the same trends as for UNED-ACCESS. The worst results are obtained for Mathematics and the best results are obtained for Biology, Micro- and Macro-economics, Computer Science, World History and Psychology. Also equally, the best models perform best in all subjects and the best performing models in average, GPT-4o and Claude-3, pass all subjects except for Mathematics.

### v2

# Eva Sánchez Salido, Roser Morante, Julio Gonzalo, Guillermo Marco Jorge Carrillo-de-Albornoz, Laura Plaza, Enrique Amigó, Andrés Fernández Alejandro Benito-Santos, Adrián Ghajari Espinosa and Victor Fresno

@@ Abstract
In this article we present UNED-ACCESS 2024, a bilingual dataset that consists of 1003 multiple-choice questions of university entrance level exams in Spanish and English. Questions are originally formulated in Spanish and manually translated into English, and have not ever been publicly released, ensuring minimal contamination when evaluating Large Language Models with this dataset. A selection of current open-source and proprietary models are evaluated in a uniform zero-shot experimental setting both on the UNED-ACCESS 2024 dataset and on an equivalent subset of MMLU questions. Results show that (i) Smaller models not only perform worse than the largest models, but also degrade faster in Spanish than in English. The performance gap between both languages is negligible for the best models, but grows up to 37% for smaller models; (ii) Model ranking on UNED-ACCESS 2024 is almost identical (0.98 Pearson correlation) to the one obtained with MMLU (a similar, but publicly available benchmark), suggesting that contamination affects similarly to all models, and (iii) As in publicly available datasets, reasoning questions in UNED-ACCESS are more challenging for models of all sizes.
@@
Eva Sánchez Salido, Roser Morante, Julio Gonzalo, Guillermo Marco Jorge Carrillo-de-Albornoz, Laura Plaza, Enrique Amigó, Andrés Fernández Alejandro Benito-Santos, Adrián Ghajari Espinosa and Victor Fresno UNED Research Center in Natural Language Processing and Information Retrievalthanks: nlp.uned.es ETSI Informática, UNED - Juan del Rosal, 16 28040 Madrid, Spain Correspondence: julio@lsi.uned.es
@@ 1 Introduction
With the recent progress in broadening the generalisation capabilities of Large Language Models (LLMs), much current research focuses on understanding their capabilities and limitations. Evaluation of generative models, such as Llama (Touvron et al., 2023b), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024a), Gemini (Anil et al., 2024), Gemma (Mesnard et al., 2024), GPT-3.5 (Brown et al., 2020), GPT-4 and GPT-4o  (Achiam et al., 2024), attempt at measuring their world knowledge, memorization and inference capabilities with benchmarks designed for these purposes. Although many benchmarks have been proposed as single-task evaluations, with the emergence of general language models such as BERT (Devlin et al., 2019), the development of more comprehensive benchmarks to measure the general capabilities of these models in a multi-task setting became popular (GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2020), Big-Bench (Srivastava et al., 2022), Big-Bench Hard (Suzgun et al., 2022), HELM (Liang et al., 2023)). A lot of emphasis is now put on assessments focusing on human-level cognitive tasks in real-world evaluation scenarios, such as exams. In this context, multiple-choice questions have emerged as one of the preferred methods for evaluating new generative models with datasets such as RACE (Lai et al., 2017), MMLU (Hendrycks et al., 2021a) or AGIEval (Zhong et al., 2023), which includes university entrance exams, law school entrance exams, mathematics competitions and lawyer qualification tests.
In this paper we present a new bilingual dataset of private exam questions, UNED-ACCESS 2024. The dataset contains 1003 multiple-choice questions in Spanish from various subjects of the UNED Access Course for Over-25s in Spanish, together with high-quality English professional translations. The dataset has been compiled at Universidad Nacional de Educación a Distancia (UNED), in the framework of the ODESIA project (Espacio de Observación de Inteligencia Artificial en Español). This dataset is complementary to the ODESIA Benchmark, a collection of comparable datasets in English and Spanish to evaluate multiple NLP discriminative tasks. The ODESIA Benchmark is used in the ODESIA Leaderboard, [^1] https://leaderboard.odesia.uned.es with the aim of measuring the effectiveness gap between Spanish and English LLMs. Two characteristics make this dataset unique: first, the contamination level is minimal, because UNED typically does not release the answers to the exam questions, which are only accessible to the teachers of each course. Second, this is a high-quality bilingual dataset, with original questions in Spanish translated into English manually by a professional translators who did not use any external software.
@@
Our main contributions are: (i) The creation of a new, private benchmark with minimal contamination for the evaluation of generative LLMs on entrance university exams in Spanish (native) and English (translated); (ii) A systematic evaluation of LLMs with the same prompts and hyperparameters for all models, in English and Spanish; (iii) Empirical evidence that the effectiveness gap of LLMs between English and Spanish is inversely correlated with model quality and size: smaller models have lower performance values in both languages and also a larger gap between their effectiveness in English and Spanish, and (iv) Empirical evidence that the relative performance of models on UNED-ACCESS 2024 is almost identical to their performance on an equivalent MMLU dataset and that results correlate with the public MMLU results, which may imply that contamination affects similarly to all language models, and that a small, sufficiently diverse dataset is representative enough to measure model performance.
@@ 2 Related work
Standard question/answer datasets include tasks such as common sense reasoning: HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), ARC (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), CommonsenseQA (Talmor et al., 2019), PIQA (Bisk et al., 2019), SIQA (Sap et al., 2019); world knowledge: TriviaQA (Joshi et al., 2017), NaturalQuestions (Kwiatkowski et al., 2019); reading comprehension: RACE (Lai et al., 2017), QuAC (Choi et al., 2018) and SQuAC (Rajpurkar et al., 2016); mathematical reasoning: MATH (Hendrycks et al., 2021b), GSM8K (Cobbe et al., 2021)); and code generation: HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021). Aggregated benchmarks like MMLU (Hendrycks et al., 2021a), AGIEval (Zhong et al., 2023), Big-Bench (Srivastava et al., 2022), and Big-Bench Hard (Suzgun et al., 2022) are also commonly used. In particular, since the introduction of GPT-3, evaluations that comprise a variety of datasets have been standardised (see for instance evaluations of GPT-3 (Brown et al., 2020), GPT-4 (Achiam et al., 2024), Llama and Llama-2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023) or Claude-3 (Anthropic, 2024)). For Spanish and its official varieties, a new benchmark for the evaluation of open generative LLMs, named La Leaderboard, has recently been released Grandury et al. (2024).
@@
Recent research has increasingly focused on multilingual evaluations of models like ChatGPT. For instance, Lai et al. (2023) evaluated ChatGPT on 37 languages ranging from high to extremely low resources, across seven tasks in a zero-shot learning setting. Their findings, along with Bang et al. (2023), show that models struggle to generalize in low-resource languages. Similarly, M3Exam (Zhang et al., 2023) assesses models like GPT-4 across nine languages, revealing persistent difficulties in handling low-resource and non-Latin script languages. GPT-4’s technical report itself extends its evaluation to a multilingual version of MMLU self-translated via Azure Translate (Achiam et al., 2024), and a professional manual translation of MMLU in 14 languages (MMMLU) has recently been released OpenAI (2024).
@@
MMMLU (OpenAI, 2024) English, Spanish and other 13 languages English Yes Yes Professional human translations Public
@@ 4 Experimental setup
Experiments were conducted using 12 generative models: four proprietary models (GPT-4-Turbo (Achiam et al., 2024), GPT-4o, GPT-3.5-Turbo (Brown et al., 2020) and Claude-3-Opus (Anthropic, 2024)) and eight open-weights models: Llama-2-7B (Touvron et al., 2023a), Llama-3-8B (Meta, 2024), Llama-3-70B, Gemma-7B (Mesnard et al., 2024), Gemma-2-27B, Mixtral-8x7B (Jiang et al., 2024a), Mistral-7B (Jiang et al., 2023) and Leniachat-Gemma-2B, all trained for instruction following. In Appendix B.1 details are provided about how the models were accessed.
We decided to apply a uniform zero-shot setting because it more closely resembles the real-world scenario in which people interact with LLMs, and it also results in a simpler and more replicable experimentation. As for the hyperparameters, the temperature was set to 0 in all models to ensure deterministic responses, minimizing creativity and focusing on factual or reasoning-based answers, which suits the dataset. The questions were provided one at a time and the prompt is fixed, providing a system prompt, a user prompt and an assistant prompt. As extra information, the prompt provided the name of the subject to which the question belongs. The language of the prompt was always in the same language of the dataset used to evaluate the model (English or Spanish) as in Zhang et al. (2023) and Achiam et al. (2024). The prompts are the following for Spanish (ES) and English (EN):
@@
In the literature, evaluation is typically based on Accuracy, the proportion of correct answers ($C$) over the total responses ($N$). However, to allow for direct comparisons across subjects with varying numbers of answer choices, we use Cohen’s Kappa coefficient, which accounts for the number of possible choices ($M$) in multiple-choice questions:
@@ 5 Results
Llama-3-70B-Instruct open 0.65 0.67 -3.44
Gemma-2-27B-Instruct open 0.64 0.66 -3.40
@@
Mixtral-8x7B-Instruct open 0.56 0.57 -2.32
Llama-3-8B-Instruct open 0.51 0.50 2.73
Mistral-7B-Instruct open 0.46 0.43 5.67
Gemma-7B-It open 0.41 0.38 8.93
Llama-2-7B-Chat open 0.32 0.25 26.04
Leniachat-Gemma-2B open 0.15 0.11 37.27
@@
Table 3 shows the average Cohen’s Kappa results by model in Spanish and English. Additional detailed results per language, model and subject are provided in Table 6 in Appendix C.1. The best overall results for Spanish are achieved by the proprietary models Claude-3-Opus, GPT-4-Turbo and GPT-4o, followed by the open models Llama-3-70B, Gemma-2-27B and Mixtral-8x7B, which outperform GPT-3.5-Turbo. In the pass margin (around 0.50 points) there is Llama-3-8B, and below are Mistral-7B, Gemma-7B, Llama-2-7B and Leniachat-2B. For English, the ranking of the models by average performance is almost identical to that of Spanish, except for an exchange in the positions of GPT-3.5 and Mixtral (GPT-3.5 performs better than Mixtral in English but worse in Spanish). At the quantitative level, the results are also very similar in English and Spanish.
@@
The last column in Table 3 shows the performance gap between English and Spanish. We calculate the percentage effectiveness gap between languages from the averaged results as:
@@
Five out of twelve models perform slightly better in Spanish than English (namely Claude-3, GPT-4, Llama-3-70B, Gemma-2-27B and Mixtral-8x7B, where the gap is negative), two of them proprietary (which perform the best in Spanish overall) and three open (also the three best models among the open-weights ones). This does not imply that these models are better in Spanish: English questions are manual translations of the original Spanish ones, and there are at least three language-related factors that affect overall performance: (i) even with high-quality manual translations, there might be translation artifacts (e.g. unusual translation choices) that make the English text harder to process; (ii) contamination: even if the questions are private, the chance that some questions have been seen by the models in Spanish is higher than in English; (iii) language competence of the models in English vs Spanish. Presumably, in the models that perform better in Spanish, effects (i) and (ii) out-weight pure language competence.
Among the proprietary models, three of them have similar performance differences (in absolute value), whilst that of GPT-3.5 is the most pronounced, in favour of English. As for the open models, the two best performing ones (Llama-3-70B and Gemma-2) have a very similar performance difference, both in favour of Spanish (and larger that the ones observed for the best proprietary models). Mixtral is the next model in the ranking and still favours Spanish, while all the other models, which do not pass (or are at the limit), favour English. In this set of models the performance difference is very pronounced, specially for Llama-2 and Leniachat, which are the worst performing models. In summary, for the best models we observe a slight advantage for Spanish, while in the worst models there is a pronounced advantage towards English. The best models are in turn larger, and therefore more likely to have been trained with more Spanish data, and more data from the web in general, whereas smaller models may rely more solely on English. However, in principle this is not the case of Leniachat, which has been conceived specifically for Spanish. The size of the models is crucial: although we do not know for certain the size of the proprietary models, in general model size is the best performance predictor, with one exception: the very similar performance between Llama-3-70B and Gemma-2-27B, with such a large difference in size, is noteworthy.
There are two cases where the gap is extremely pronounced (in favour of English), coinciding with the two worst-performing models: Leniachat, which is the smallest model, and Llama-2, which is the oldest of the evaluated models.
There is a strong, statistically significant inverse correlation between model quality and language gap: Pearson’s correlation between English performance and EN-ES gap is -0.87 with a p-value of 0.0002; and Pearson’s correlation between performance in Spanish and language gap is -0.89 (p=9.83e-05). The performance of less powerful models degrades faster in Spanish than in English.
@@
A direct comparison between the public results on MMLU (Hendrycks et al., 2021a) and UNED-ACCESS 2024 is not adequate: the dataset sizes are different (MMLU has more than 14000 test instances) and the subjects and levels differ: MMLU ranges in difficulty from an elementary level to an advanced professional level, whereas UNED-ACCESS 2024 covers only university-entrance level. Also, most public evaluations of MMLU are 5-shot and even include chain of thought prompting, as compared to the zero-shot setting of our experiments. For these reasons we have performed our own evaluation of the models, using the same methodology described earlier, and selecting the subset of 2822 MMLU questions (MMLU-10) from ten topics which correspond most closely to the level and topics of UNED-ACCESS: high school biology, high school chemistry, high school computer science, elementary mathematics, high school mathematics, high school physics, high school macroeconomics, high school microeconomics, high school psychology and high school world history.
MMLU from publications MMLU-10 from zero-shot experiments
GPT-4o 88.7 83.40
Claude-3-Opus 86.8 (5-shot) 83.71
GPT-4-Turbo 86.5 81.61
Llama-3-70B-Instruct 82.0 Llama-3-70B (5-shot) 75.47
Gemma-2-27B-Instruct 75.2 Gemma-2 (5-shot) 76.61
Mixtral-8x7B-Instruct 70.6 Mixtral 8x7B (5-shot) 65.34
GPT-3.5-Turbo 70.0 63.74
Llama-3-8B-Instruct 68.4 Llama-3-8B (5-shot) 61.54
Mistral-7B-Instruct 58.4 (5-shot) 54.24
Gemma-7B-It 53.3 (5-shot) 42.50
Llama-2-7B-Chat 45.3 Llama-2-7B (5-shot) 35.08
Leniachat-Gemma-2B - 29.81
First, we have checked whether the results that we obtain on MMLU-10 with our methodology correlate with the published results. Table 4 provides the MMLU results published in the official benchmark [^3] https://paperswithcode.com/sota/multi-task-language-understanding-on-mmlu (or, in their absence, in the respective model reports) and the results of applying our methodology on MMLU-10. Both evaluations are reported in terms of Accuracy, rather than Cohen’s Kappa, because published results are always presented with this metric. This is the only table in this paper that presents results using Accuracy. Pearson’s correlation between these two sets of results is 0.9863 with a p-value of 2.31e-08, which means that results are highly correlated. This high correlation validates our experimentation, as it implies that our choices of prompts and hyperparameters, and zero-shot setting, yield results consistent with those found in the literature.
This analysis allows us to compare the results obtained on MMLU-10 with the results on UNED-ACCESS 2024 in English and check whether they correlate. Table 5 shows both results in terms of Cohen’s Kappa (recall that the MMLU questions all have 4 answers). Pearson’s correlation between these results is 0.98 (p=5.04e-08): both results have almost perfect, statistically significant correlation. This consistency reinforces the reliability of results obtained with both datasets.
@@
The fact that the results on both datasets are so similar, in spite of the fact that UNED-ACCESS should have minimal contamination, implies that either the questions in UNED-ACCESS are somewhat easier in average, or that there is some kind of indirect contamination: some questions might be recurring for some subjects, and models might have seen variations of the questions even if they do not have access to the original ones. Also, the similarity of the results seems to indicate that the translation effect is not decisive.
In Table 7 of Appendix C.2 we also provide the evaluation results on the MMLU-10 dataset per subject, which shows the same trends as for UNED-ACCESS 2024. The worst results are obtained for Mathematics and the best results are obtained for Biology, Micro- and Macro-economics, Computer Science, World History and Psychology. Also equally, the best models perform best in all subjects and the best performing models in average, GPT-4o and Claude-3, pass all subjects except for Mathematics.